What gauge pressure must a pump produce in order to pump water from the bottom of

Grand Canyon (730 m elevation) in Pascals and in atmosphere?


1
Expert's answer
2021-06-17T15:07:39-0400

By the definition of the gauge pressure we have:


"P_{gauge}=P_{abs}-P_{atm}=\rho_{water}g\Delta h,""P_{gauge}=\rho_{water}g(h_2-h_1)."

From the definition of the question, we know "h_1=730\ m". But, we don't know "h_2". Let's suppose that "h_2=1200\ m". Then, we get:


"P_{gauge}=1000\ \dfrac{kg}{m^3}\cdot9.8\ \dfrac{m}{s^2}\cdot(1200\ m-730\ m)=4.61\cdot10^6\ Pa,""P_{gauge}=4.61\cdot10^6\ Pa\cdot\dfrac{1\ atm}{101325\ Pa}=45.5\ atm."